# Cone structures from a dynamical and probabilistic viewpoint

### v1

@@ 1. Introduction
Moreover, by projecting the original Reeb field on the $5$-manifold to $S^{3}$, the author constructed a cone structure, i.e. a choice of conical set $C=\{C_{p}\}_{p\in S^{3}}$ depending smoothly on $p$. Here, a conical set is a subset of the tangent bundle to $M$, invariant under the natural $\mathbb{R}^{+}$ scaling action. The cone structure is trivial at a point $p$ if $C_{p}=\{0\}$. The cone structure $C$ which arises on $S^{3}$ is $2$-dimensional, i.e. it is the image of a map with a two-dimensional domain. The orbits of the original flow project, by construction, to smooth curves tangent to the cone structure. With the orientation given, the tangent plane $TP$ to the page determines a half-space of positive directions. We will call the intersection of an affine plane parallel to $TP$, with the cone, a base of the cone, see e.g. [5]. There is a natural Riemannian metric on $S^{3}$ associated to a suitable almost complex structure and the averaged contact form, which will be implicitly used in what follows to measure angles and distances. In this context, we have the following.
@@ 3. Invariants for cone structures
is the contact volume of the ambient manifold. If $\tau_{n}(\Gamma):\mathcal{P}_{C}\rightarrow\mathbb{R}$ represents the $n$-th return time function (i.e. the time it takes for the orbit to return $n$ times to $P$), with respect to an arbitrary section $\Gamma$, it would be interesting to provide information on the growth rate of the limit
@@
$E(\Gamma)=\int_{P}\frac{x}{\mu(P)}d\mu(x)=1=\int_{P}xd\mu(x),$
is the center of mass of the page $P$, and
@@ 4. Brownian motion for cone structures
This type of process can also be adapted to an open book with an adapted cone structure (e.g. for pur running example on $S^{3}$), by imposing that the Brownian motion is interior to the cone structure, and has a positive drift in the open book angle direction. In this situation, given a section $\Gamma$ for $\mathcal{P}_{C}$, the function $\tau(\Gamma)$ can be interpreted as the average return time to a fixed page, of a curve in the section. From standard results of Brownian motion, we obtain the following.
@@
Note that in a deterministic situation (e.g. coming from a Hamiltonian system), the above proposition is compatible with Poincaré recurrence. We will then call the above, probabilistic Poincaé recurrence. Heuristically, from the perspective of quantum mechanics, and in particular, the Feynmann path integral, one can interpret the smooth curves in $\mathcal{P}_{C}$ as the “classical solutions”, and their probabilistic version (the instantiations of the process $V_{t}$) as the “quantum oscillations around the classical solution”.

### v3

@@ 1. Introduction
Moreover, by projecting the original Reeb field on the $5$-manifold to $S^{3}$, the author constructed a cone structure, i.e. a choice of conical set $C=\{C_{p}\}_{p\in S^{3}}$ depending smoothly on $p$. Here, a conical set is a subset of the tangent bundle to $M$, invariant under the natural $\mathbb{R}^{+}$ scaling action. The cone structure is trivial at a point $p$ if $C_{p}=\{0\}$. The cone structure $C$ which arises on $S^{3}$ is $2$-dimensional, i.e. it is the image of a map with a two-dimensional domain. The orbits of the original flow project, by construction, to smooth curves tangent to the cone structure. With the orientation given, the tangent plane $TP$ to the page determines a half-space of positive directions. A base for a cone inside a tangent space is a section of the cone, i.e. a continuous choice of vector in each direction tangent to the cone; a base for a cone structure is a continously varying choice of bases for each tangent space, see e.g. [5]. A based cone structure is a cone structure for which a continuously varying base has been chosen. Such a based cone structure naturally arises in the picture for the circular, restricted three-body problem. There is a natural Riemannian metric on $S^{3}$ associated to a suitable almost complex structure and the averaged contact form, which will be implicitly used in what follows to measure angles and distances. In this context, we have the following.
@@ 3. Invariants for cone structures
is the contact volume of the ambient manifold. In general, $\mbox{CAL}_{\Gamma}(A)$ agrees with the volume of the suspension of $A$.
If $\tau_{n}(\Gamma):\mathcal{P}_{C}\rightarrow\mathbb{R}$ represents the $n$-th return time function (i.e. the time it takes for the orbit to return $n$ times to $P$), with respect to an arbitrary section $\Gamma$, it would be interesting to provide information on the growth rate of the limit
@@
$E(\Gamma)=\int_{P}\frac{x}{\mu(P)}d\mu(x)=\int_{P}xd\mu(x),$
is the center of mass of the page $P$ (i.e. the “origin” of the disk), and
@@ 4. Brownian motion for cone structures
This type of process can also be adapted to an open book with an adapted cone structure (e.g. for our running example on $S^{3}$), by imposing that the Brownian motion is interior to the cone structure, and has a positive drift in the open book angle direction. In this situation, given a section $\Gamma$ for $\mathcal{P}_{C}$, the function $\tau(\Gamma)$ can be interpreted as the average return time to a fixed page, of a curve in the section. From standard results of Brownian motion, we obtain the following.
@@
Note that in a deterministic situation (e.g. coming from a Hamiltonian system), the above proposition is compatible with Poincaré recurrence. We will then call the above, probabilistic Poincaré recurrence. Heuristically, from the perspective of quantum mechanics, and in particular, the Feynmann path integral, one can interpret the smooth curves in $\mathcal{P}_{C}$ as the “classical solutions”, and their probabilistic version (the instantiations of the process $V_{t}$) as the “quantum oscillations around the classical solution”.